OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10)

OpenStudy (anonymous):

A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

OpenStudy (anonymous):

B and D are the same? Either way y = x + 6 is parallel to y = x + 8.